Find the centroid of the region bounded by the curves y= x3 and y=4x in the fourth quadrant. Sketch the bounded region.


1
Expert's answer
2021-06-14T17:53:58-0400

Here the region is bounded in first and third quadrants.


So, there is no centroid at 4th quadrant.


Centroid at 1st quadrant is,


Mx = (1/2)∫ab (f(x)2 - g(x)2) dx


= (1/2)((16x3/3) - (x7/7))01

= 2.5923


My = ∫ab x(f(x)-g(x)) dx


= ((4x3/3)-(x5/5))"_0^1"

= 3.85714


M = ∫ab (f(x)-g(x)) dx


= (2x2 - x4/4)"_0^1"

= 1.75


Thus, Centroid,(xi,yi) = (My/M,Mx/M) = (2.20408,1.48289)